OpenStudy (anonymous):

find f'(x) f(x)=x/(1+x) using f(x+h)-f(x) / h

myininaya (myininaya):

\[f(x+h)=\frac{x+h}{1+x+h}\]

myininaya (myininaya):

\[\lim_{h \rightarrow 0}\frac{\frac{x+h}{1+x+h}-\frac{x}{1+x}}{h}\]

myininaya (myininaya):

\[\lim_{h \rightarrow 0}\frac{(1+x+h)(1+x)}{(1+x+h)(1+x)} \cdot \frac{\frac{x+h}{1+x+h}-\frac{x}{1+x}}{h}\]

myininaya (myininaya):

\[\lim_{h \rightarrow 0}\frac{(x+h)(1+x)-x(1+x+h)}{(1+x+h)(1+x)h}\]

myininaya (myininaya):

\[\lim_{h \rightarrow 0}\frac{x+x^2+h+xh-x-x^2-xh}{(1+x+h)(1+x)h}\]

myininaya (myininaya):

\[\lim_{h \rightarrow 0}\frac{h}{(1+x+h)(1+x)h}=\lim_{h \rightarrow 0}\frac{1}{(1+x+h)(1+x)}\]

myininaya (myininaya):

\[\frac{1}{(1+x+0)(1+x)}=\frac{1}{(1+x)^2}\]

OpenStudy (anonymous):

wow thx myininaya!

OpenStudy (anonymous):

i don't get it could you write it again?

myininaya (myininaya):

don't mind satellite hes always out to punish me

OpenStudy (anonymous):

lol